# Uncovering Cultural Representation Disparities in Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities across a range of tasks, yet concerns about their potential biases exist. This work investigates the extent to which prominent VLMs exhibit cultural biases by evaluating their performance on an image-based country identification task at a country level. Utilising the geographically diverse Country211 dataset, we probe several large vision language models (VLMs) under various prompting strategies: open-ended questions, multiple-choice questions (MCQs), including challenging setups like multilingual and adversarial settings. Our analysis aims to uncover disparities in model accuracy across different countries and question formats, providing insights into how training data distribution and evaluation methodologies might influence cultural biases in VLMs. The findings highlight significant variations in performance, suggesting that while VLMs possess considerable visual understanding, they inherit biases from their pre-training data and scale that impact their ability to generalize uniformly across diverse global contexts.
@@ 1 Introduction
Vision-Language Models (VLMs) have rapidly advanced, demonstrating exceptional capabilities in integrating visual and textual information for a wide array of tasks, from image captioning to visual question answering (Liu et al., 2024; Alayrac et al., 2022; Wang et al., 2024). These models are increasingly being deployed in diverse applications, impacting areas such as education, healthcare, and public services globally (Zhang et al., 2024).
@@
In order to quantify cultural disparities in VLMs, we adopt image-based country identification as a concrete proxy task in which a model must infer an image’s country of origin solely from visual cues, while also providing a justification. Prior work has shown that geolocation tasks reveal representational imbalances in visual models, as performance often correlates with the prevalence of training data from different regions (Pouget et al., 2024). Using a dataset, with a balanced distribution of images from each country, we ensure that the observed accuracy gaps stem from model bias rather than dataset imbalance. Further, the complex nature of images used could measure VLMs’ ability to distinguish similar cultures/countries.
@@
    We systematically probe VLMs under varied settings—open-ended and multiple-choice questions (MCQs) with both random and culturally similar distractors—alongside multilingual prompts in five languages, to capture nuanced cultural and linguistic disparities.
@@
    Our findings show that VLM biases do not consistently favour Western countries; instead, biases often reflect overrepresentation of certain popular countries (e.g., India, USA) in the training data, suggesting a more complex bias landscape.
@@ 2 Related Works
Prior Work Eval Method Multilingual? Adversarial? Categories Total Sample Count Domain
@@ 3 Dataset Used
Existing benchmarks highlight cultural blind spots in VLMs, but they generally either cover fewer categories or countries or are restricted to specialized domains. Our work differs by using an image-based country-identification task over 211 countries, providing much broader geographic coverage and adversarial probing. Further, the datasets utilized in the prior works utilize a images that might be easier to classify, including but not limited to close up shots of food items, popular monuments being the primary object in an image etc. The dataset we utilized introduces a lot of noise and randomness in a majority of images as seen in Figure 2. For instance, the examples from Norway, India and Egypt might be easy to classify, but the examples from Afghanistan and Kuwait require grasping certain features and their associated knowledge i.e the headgear pattern of the Kuwait image and how it is different from other countries in the region. Th example from Afghanistan requires noticing the afghan flag, while the appearance of the person in the left may try to mislead the VLM due to an appearance of different ethnicity.
@@ 4 Experiments
    multiple-choice questions (with random distractors)
@@
    multiple-choice questions (with similar distractors)
Image perturbations:The open-ended experiments were re-done with following adversarial changes:
@@ 5 Results
The results for experimental setting over countries of each region can be seen in Table 2.
MCQA
Region Open-Ended Similar Random
@@ 5.2 Effect of Image Perturbations on Results
Figure 5: Average model confidence, given the original images (left), grayscale images (middle), and rotated images (right). GPT4o, GeminiFlash, and Gemma27B are most sensitive to image perturbations.
@@ 5.4 Image Feature categories VS accuracy
Apart from the experiments, the original 21.1 K images were also labeled multi-way based on the key features they contain using larger VLMs like Gemini-2.5-Pro, o4-mini, Grok-2-Vision. Later a majority vote of each label was considered. The quality was later manually verified over a subset by multiple people. We have used 9 sub-categories for this categorization. The descriptions of each of these categories can be seen in Table 3. A large variance was observed between each feature category and the country level accuracies obtained. Additionally there was also a large variation between how accuracy was affected for each country/feature based on model/perturbation used. This can be also be seen in Figure 13. The extent to which each category’s images were recognized by VLMs can be seen in Figure 4. External architecture and native language texts’ presence in the background helped the VLMs recognize the culture better compared to the other features.
@@ 5.5 Distribution of Predicted countries
Notably, all models consistently overpredict certain countries—particularly the USA, India, and Brazil—regardless of actual ground truth. We hypothize that these countries are likely overrepresented in the models’ pretraining data or benefit from more visually distinctive cues. Biases seem to cluster around a few highly represented or visually salient countries rather than reflecting broader geopolitical landscape.
These results show that model predictions are likely highly influenced by data availability and image characteristics rather than a generic global bias. It also underscores the need for better interpretabilty regarding the geographic composition of VLM training datasets to fully understand such biases.
@@ 6 Discussion
Our study presents a comprehensive analysis of cultural biases in Vision-Language Models (VLMs) using a geographically balanced dataset across 211 countries. The key difference between the samples in our datasets and those from the other works is the complexity of images provided. In most of the images, we used it is difficult to understand and predict the country of origin without a strong vision ability. For example, the Afghanistan example from Figure 2 is difficult to classify without spotting the hint of Afghan flag from one of the people from the image.
We evaluated popular models across multiple prompting strategies, e.g. open-ended, multiple-choice (random and similar distractors), and multilingual settings. Open-ended formats showed the greatest disparity in country-level accuracy, particularly in underrepresented regions such as Central Africa and parts of South America.The use of culturally similar distractors proved to be the most effective in revealing fine-grained errors, highlighting limitations in models’ cultural discrimination abilities.
We further assessed the models’ robustness to image perturbations like gray-scaling and rotation. While gray-scaling affected only a few specific countries, rotation led to a broad and uniform drop in performance, confirming that VLMs rely heavily on image orientation. We further observed that performance also varied by semantic image content—categories like architecture, textual cues, and attire were more predictive of cultural origin, especially in unaltered images.
Language variation in prompts had minimal impact on average accuracy, though countries closely tied to the input language (e.g., Spain with Spanish, Brazil with Portuguese) showed slight gains. However, this trend was inconsistent and did not generalize across all culturally linked regions.
Finally, our misclassification analysis shows that models frequently confuse images from low-resource or visually ambiguous countries with a few dominant nations, reinforcing the role of training data bias. These findings emphasize the need for transparent reporting of dataset composition in VLM development and call for more robust, culturally diverse benchmarks to ensure equitable global performance
@@ 7 Conclusion
We investigated cultural biases in Vision-Language Models using a diverse country-level image classification task. Our findings show that biases are not uniformly Western but instead reflect overrepresentation of certain countries in training data. Model performance varied across prompt types, languages, image features, and perturbations—highlighting limitations in robustness and cultural generalization.
These results call for greater transparency in dataset composition and the need for more culturally inclusive evaluation methods to ensure fairer and more globally representative VLMs.
@@ Appendix G Country wise accuracies in each experimental setting
Country name Open-Ended MCQs with MCQs with
Similar choices Random choices
@@
Country name Open-Ended MCQs with MCQs with
Similar choices Random choices
@@
Country name Open-Ended MCQs with MCQs with
Similar choices Random choices
@@
Country name Open-Ended MCQs with MCQs with
Similar choices Random choices
@@
Country name Open-Ended MCQs with MCQs with
Similar choices Random choices

### v3

@@ Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities across a range of tasks, yet concerns about their potential biases exist. This work investigates the extent to which prominent VLMs exhibit cultural biases by evaluating their performance on an image-based country identification task at a country level. Utilising the geographically diverse Country211 (OpenAI, 2021) dataset, we probe several VLMs under various prompting strategies: open-ended questions, multiple-choice questions (MCQs), including challenging setups like multilingual and adversarial settings. Our analysis aims to uncover disparities in model accuracy across different countries and question formats, providing insights into how training data distribution and evaluation methodologies might influence cultural biases in VLMs. The findings highlight significant variations in performance, suggesting that while VLMs possess considerable visual understanding, they inherit biases from their pre-training data and scale that impact their ability to generalize uniformly across diverse global contexts.
@@ 1 Introduction
VLMs have rapidly advanced, demonstrating exceptional capabilities in integrating visual and textual information for a wide array of tasks, from image captioning to visual question answering (Liu et al., 2024; Alayrac et al., 2022; Wang et al., 2024). These models are increasingly being deployed in diverse applications, impacting areas such as education, healthcare, and public services globally (Zhang et al., 2024).
@@
In order to quantify cultural disparities in VLMs, we adopt image-based country identification as a concrete proxy task in which a model must infer an image’s country of origin solely from visual cues, while also providing a justification. Prior work has shown that geolocation tasks reveal representational imbalances in visual models, as performance often correlates with the prevalence of training data from different regions (Pouget et al., 2024).
@@
    We systematically probe VLMs under varied settings—open-ended and multiple-choice questions (MCQs) with both random and culturally similar distractors—alongside multilingual prompts in five languages, to capture nuanced cultural and linguistic disparities.  [^1] Due to cultural similarities, misclassification among similar countries is more likely than misclassification with an unrelated country. MCQ with random and similar distractors tested the VLMs in both scenarios as to whether misclassification would occur when all distractors are neither neighboring nor similar countries
@@
    Our findings show that VLM biases do not consistently favour Western countries; instead, biases often reflect over representation of certain popular countries (e.g., India, USA) in the training data  [^2] For deliberately under specified inputs without country names, the generated images most reflect the surroundings of the United States followed by India. (Basu et al., 2023), suggesting a more complex bias landscape.
@@ 2 Related Works
[…]
@@
**The vales in brackets indicate the features in the Country recognition task subset, while the first values indicate that of the whole dataset.
@@ 3 Dataset Used
Existing benchmarks highlight cultural blind spots in VLMs, but they generally either cover fewer categories or countries or are restricted to specialized domains. Our work differs by using an image-based country-identification task over 211 countries, providing much broader geographic coverage and adversarial probing. Further, the datasets utilized in the prior works utilize a images that might be easier to classify, including but not limited to close up shots of food items, popular monuments being the primary object in an image etc. The dataset we utilized introduces a lot of noise and randomness in a majority of images as seen in Figure 2444The images were part of OpenAI’s YFCC100M and come with pre-verified country labels. Although some samples might be difficult to classify even for a native, The primary goal was to uncover cultural biases using the features the VLMs could probably misclassify it with a culturally similar or neighboring country, but frequently misclassify it with a very dissimilar country.. For instance, the examples shown from UK, India and Egypt might be easy to classify, but the examples from Afghanistan and Kuwait require grasping certain features and their associated knowledge i.e the headgear pattern of the Kuwait image and how it is different from other countries in the region. Th example from Afghanistan requires noticing the afghan flag, while the appearance of the person in the left may try to mislead the VLM due to an appearance of different ethnicity.
@@ 4 Experiments
    MCQs (with random distractors)
@@
    MCQs (with similar distractors)
Image perturbations: Open-ended experiments were done with these adversarial changes:
@@ 5 Results
The results for experimental setting over countries of each region can be seen in Table 3. Additionally Table 2 demonstrates the statistical significance of the results which presents Pearson’s $\chi^{2}$ and Cochran’s Q test results for every evaluation condition and its subcategories. The $\chi^{2}$ p-values of 0.0 (effectively under flowing) reveal that the six VLMs differ in accuracy in a highly significant way, confirming that VLM biases drive overall performance differences. In contrast, the Cochran’s Q p-values of 1.0 indicate no significant variation among sub-conditions (whether comparing original, rotated, or grayscale images, or across the five languages), showing that these perturbations do not meaningfully alter each VLM’s overall accuracy, but do cause changes in country wise accuracy as seen in Figure 14 and Figure 15.
[…]
Open-ended
Original 5851.81 0.0$\sim$ 0.0∼ 0.0 -0.727 1.0$\sim$ 1.0∼ 1.0
Rotated 4666.49 0.0$\sim$ 0.0∼ 0.0 -0.952 1.0$\sim$ 1.0∼ 1.0
Greyscale 4280.08 0.0$\sim$ 0.0∼ 0.0 -0.725 1.0$\sim$ 1.0∼ 1.0
MCQ-S
Overall 26274.855 0.0$\sim$ 0.0∼ 0.0 -0.151 1.0$\sim$ 1.0∼ 1.0
ENG 5232.00 0.0$\sim$ 0.0∼ 0.0 -0.145 1.0$\sim$ 1.0∼ 1.0
HIN 5197.50 0.0$\sim$ 0.0∼ 0.0 -0.153 1.0$\sim$ 1.0∼ 1.0
POR 5309.00 0.0$\sim$ 0.0∼ 0.0 -0.155 1.0$\sim$ 1.0∼ 1.0
SPA 5256.72 0.0$\sim$ 0.0∼ 0.0 -0.151 1.0$\sim$ 1.0∼ 1.0
ZHO 5345.21 0.0$\sim$ 0.0∼ 0.0 -0.155 1.0$\sim$ 1.0∼ 1.0
MCQ-R
Overall 23855.61 0.0$\sim$ 0.0∼ 0.0 -0.074 1.0$\sim$ 1.0∼ 1.0
ENG 4835.90 0.0$\sim$ 0.0∼ 0.0 -0.073 1.0$\sim$ 1.0∼ 1.0
HIN 4584.62 0.0$\sim$ 0.0∼ 0.0 -0.073 1.0$\sim$ 1.0∼ 1.0
POR 4873.45 0.0$\sim$ 0.0∼ 0.0 -0.076 1.0$\sim$ 1.0∼ 1.0
SPA 4782.98 0.0$\sim$ 0.0∼ 0.0 -0.074 1.0$\sim$ 1.0∼ 1.0
ZHO 4801.08 0.0$\sim$ 0.0∼ 0.0 -0.076 1.0$\sim$ 1.0∼ 1.0
Table 2: Statistical tests for each evaluation condition and subcategory : Open ended, MCQ-Random (MCQ-R), MCQ-Similar (MCQ-S)
[…]
@@ 5.2 Effect of Image Perturbations on Results
Figure 5: Average model confidence, given the original images (left), grayscale images (middle), and rotated images (right). GPT4o, Gemini-Flash, and Gemma-27B are most sensitive to image perturbations.
@@ 5.4 Image Feature categories VS accuracy
Apart from the experiments, the original 21.1 K images were also labeled multi-way based on the key features they contain using larger VLMs like Gemini-2.5-Pro, o4-mini, Grok-2-Vision. Later a majority vote of each label was considered. The quality was later manually verified over a subset by multiple people  [^7] Feature category labels were verified on a subset of 10% samples equally distributed over all countries, with 2 people verifying labels, in cases with no consensus between the two, the third annotator was used.. We have used 9 sub-categories for this categorization. The descriptions of each of these categories can be seen in Table 4. A large variance was observed between each feature category and the country level accuracies obtained. Additionally there was also a large variation between how accuracy was affected for each country/feature based on model/perturbation used. This can be also be seen in Figure 13. The extent to which each category’s images were recognized by VLMs can be seen in Figure 4. External architecture and native language texts’ presence in the background helped the VLMs recognize the culture better compared to the other features.
@@ 5.5 Distribution of Predicted countries
Notably, all models consistently overpredict certain countries—particularly the USA, India, and Brazil—regardless of actual ground truth. We hypothesize that these countries are likely overrepresented in the models’ pretraining data or benefit from more visually distinctive cues. Biases seem to cluster around a few highly represented or visually salient countries rather than reflecting broader geopolitical landscape.
These results show that model predictions are likely highly influenced by data availability and image characteristics rather than a generic global bias. It also underscores the need for better interpretability regarding the geographic composition of VLM training datasets to fully understand such biases.
@@ 6 Discussion
Our study presents a comprehensive analysis of cultural biases in Vision-Language Models (VLMs) using a geographically balanced dataset across 211 countries. We evaluated popular models across multiple prompting strategies, e.g. open-ended, multiple-choice (random and similar distractors), and multilingual settings. Open-ended formats showed the greatest disparity in country-level accuracy, particularly in underrepresented regions such as Central Africa and parts of South America.The use of culturally similar distractors proved to be the most effective in revealing fine-grained errors, highlighting limitations in models’ cultural discrimination abilities.
We further assessed the models’ robustness to image perturbations like gray-scaling and rotation. While gray-scaling affected only a few specific countries, rotation led to a broad and uniform drop in performance, confirming that VLMs rely heavily on image orientation. We further observed that performance also varied by semantic image content—categories like architecture, textual cues, and attire were more predictive of cultural origin, especially in unaltered images. Language variation in prompts had minimal impact on average accuracy, though countries closely tied to the input language (e.g., Spain with Spanish, Brazil with Portuguese) showed slight gains. However, this trend was inconsistent and did not generalize across all culturally linked regions.
Finally, our misclassification analysis shows that models frequently confuse images from low-resource or visually ambiguous countries with a few dominant nations, reinforcing the role of training data bias.
@@ 7 Conclusion
Our findings show that biases are not uniformly Western but instead reflect over representation of certain countries in training data. Model performance varied across prompt types, languages, image features, and perturbations—highlighting limitations in robustness and cultural generalization. These results call for greater transparency in dataset composition and the need for more culturally inclusive evaluation methods to ensure fairer and more globally representative VLMs.
@@ Appendix G Country wise accuracies in each experimental setting
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Figure 35: Examples from ours (1st,4th) as well as other works : GIMMICK (2nd), CVQA (3rd) : The 1st and 4th image have the key features required for classifying the image accurately, occupying a tiny portion of the image making it relatively difficult i.e the flag patch in image 1 ,and name of mountain in image 4’s signboard. While, in Image 2 and image 3 , the key features i.e the text on attire or the (car, city name signboard, multilingual texts on left) make the samples relatively easier to classify